# Theory of Super-Resolution Data Assimilation with Conditional Variational Autoencoders: Using Super-Resolution Operators as Background Error Covariance Matrices

### v3

@@ VI.5 Computation time of SRDA
Lastly, we confirm that the computational time of the SRDA is much less than that of the EnKF. The elapsed times for the simulations from $t=0$ to 20 were measured using a single process on a CPU (Intel Xeon Gold 6326). For the SRDA, in addition to the assimilation by the encoder, the LR forecast was super-resolved by the SR model at every output time step ($\Delta t=0.25$). The average wall time was 21.3 s for the SRDA, considerably less than 205.7 s for the EnKF. This difference is mainly due to the calculation of the ensemble evolution, which is not required in the SRDA.
VII Conclusions
This study has proposed a theory of unsupervised SRDA using the CVAE. We have shown that a new DA framework using the ELBO encompasses the 3D-Var formalism (Section III). Since ELBO serves as the loss function for CVAEs, the assimilation can be executed using the CVAE. In this process, it is essential to use a non-diagonal background error covariance matrix to incorporate the spatial correlations of forecast states into the CVAE, similar to the 3D-Var. Instead of using such non-diagonal matrices, we have further extended the framework by using the non-local nature of SR operations to achieve DA (Section IV). In other words, SR operators can serve as background error covariance matrices. This result provides a theoretical foundation for combining SR and DA.
The combination of SR models and CVAEs enables the assimilation of HR observations into LR forecasts that are computed with LR fluid models. This approach eliminates the need for numerical integration using HR fluid models, thereby reducing computational burden. Specifically, the CVAE’s encoder estimates the HR analysis from the input of LR forecasts and HR observations, while the decoder acts as the observation operator and reconstructs the HR observations from the estimated HR analysis. The resulting HR analysis is then downsampled to the LR using algebraic methods, such as linear interpolation, serving as the initial condition for the LR fluid model in the next assimilation cycle.
@@ VII Conclusions
The effectiveness of the proposed method was evaluated through numerical experiments using an idealized barotropic ocean jet system (Sections V and VI). Compared to the EnKF, the SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times. These results suggest the potential of SRDA as an efficient approach for integrating observations into CFD simulations.
There are several future research topics, as SRDA has been recently developed. Although the proposed theory enables learning of nonlinear observation operators, the proof of concept in the present study has examined only the linear observation operator. Furthermore, advanced training methods, such as online trainingRasp (2020) and loss balancing for VAEs,Asperti and Trentin (2020) will be important to develop SRDA, whereas a simple offline training method was employed in our study (Section V.4). These points will be addressed in a more realistic experimental setup. In practical applications, data partitioning will be important because the data size for three-dimensional flows is larger, which would require the division of snapshots into small patches.Ouala et al. (2018); Yang and Grooms (2021) Non-Gaussianity is also important in more realistic DA, which has recently been addressed by VAEsXiao et al. (2024) and normalizing flows.Xie et al. (2023) The theory proposed in our study will provide a foundation for using such advanced techniques, leading to further development of SRDA.
Acknowledgements.
@@
Data Availability Statement
The data and source code that support the findings of this study are preserved and openly available at https://github.com/YukiYasuda2718/srda-cvae.
@@ Appendix A Hyperparameters of the SR model and the CVAE
We list here the hyperparameters of the SR model and the CVAE shown in Fig. 5. Table 2 lists the hyperparameters for the convolution layers in the SR model and the CVAE. The kernel size of all convolutions is 3. The slope of all leaky ReLUMaas et al. (2013) is $-0.01$. The upscale factor of all pixel shuffle operationsShi et al. (2016) is 2. The resizing in the SR model is performed by bicubic interpolation. The weight parameters of all convolution layers are randomly initialized using uniform distributions.He et al. (2015)

### v4

@@ 6.5 Computation time of SRDA
The computational time of the SRDA is much less than that of the EnKF. The elapsed times for the simulations from $t=0$ to 20 were measured using a single process on a CPU (Intel Xeon Gold 6326). For the SRDA, in addition to the assimilation by the encoder, the LR forecast was super-resolved by the SR model at every output time step ($\Delta t=0.25$). The average wall time was 21.3 s for the SRDA, considerably less than 205.7 s for the EnKF. This difference is mainly due to the calculation of the ensemble evolution, which is not required in the SRDA.
7 Conclusions
This study has proposed a theory of unsupervised SRDA with the CVAE. Using the ELBO, we have developed an unsupervised theory of DA, which is an extension of the 3D-Var (Section 3). Since ELBOs generally serve as the loss function for CVAEs, the assimilation is feasible using the CVAE. In this theory, it is essential to use a non-diagonal background error covariance matrix, as in the 3D-Var. Instead of using such non-diagonal matrices, we have extended the theory by using the non-locality of SR to perform DA (Section 4). This extension provides a theoretical foundation for combining SR and DA. The combination of SR models and CVAEs enables us to assimilate HR observations into LR forecasts computed with LR fluid models. This approach eliminates the need for time integration of HR fluid models, thereby reducing computational burden.
@@ 7 Conclusions
The effectiveness of the proposed method was evaluated via numerical experiments using an idealized barotropic ocean jet system (Sections 5 and 6). Compared to the EnKF, the SRDA demonstrated its ability to obtain accurate HR inferences within shorter computational times. Our results suggest the potential of SRDA as a promising approach for integrating observations into CFD simulations while obtaining HR inferences with a low computational cost.
There are several future research topics for SRDA. The proof of concept here has only examined the linear observation operator, although the proposed SRDA theoretically allows learning of nonlinear observation operators. Furthermore, advanced training methods, such as online training [63, 64] and loss balancing for VAEs [68], will be important to develop SRDA. As for generalizability, we have not evaluated to what extent the CVAE remains effective when the statistical properties of the input data differ. These aspects should be addressed in a more realistic experimental setup. Indeed, the fluid system used here is simple compared to actual atmospheric and oceanic systems. It is necessary to evaluate the proposed SRDA for more complex systems.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
@@ Appendix A Hyperparameters of the SR model and the CVAE
We list the hyperparameters of the SR model and the CVAE shown in Fig. 7. Table 2 shows the numbers of channels and strides in all convolution layers. The kernel size of all convolutions is 3. The slope of all leaky ReLU [59] is $-0.01$. The upscale factor of all pixel shuffle operations [57] is 2. The resizing in the SR model is performed by bicubic interpolation. The weight parameters of all convolution layers are initialized randomly using uniform distributions [80].
@@
Data availability
The data and source code that support the findings of this study are preserved and openly available at our Zenodo repository (https://doi.org/10.5281/zenodo.13958042). These files are also available at our GitHub repository (https://github.com/YukiYasuda2718/srda-cvae/releases/tag/v0.2.1).
Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used Claude 3.5 only for English editing. After using this service, the authors reviewed and edited the content as needed. The authors take full responsibility for the content of the publication.